The weight of a liquid varies directly as its volume V. If the weight of the liquid in a cubical container 5 cm on a side is 375g, find the weight of the liquid in a cubical container 4 cm on a side.
Answer by Alan3354(69443)

The weight of a liquid varies directly as its volume V. If the weight of the liquid in a cubical container 5 cm on a side is 375g, find the weight of the liquid in a cubical container 4 cm on a side.
-------------
Vol = s^3 = 5^3 = 125 cc
Density = 375/125 = 3 gm/cc
===============
Multiply the volume of the 2nd cube by the density.